1. Water at 25.0C exposed to air at 1.50 atm pressure can contain a maximum of 13.7 mg/L of nitrogen gas. Calculate the value of Henry’s law constant (mol/L  atm) for nitrogen at this temperature. The mole fraction of nitrogen in air is 0.78084.

2. If 125 mL of aqueous solution with a density of 0.810 g/mL contains 10.7 mL of CH3CH2OH (ethanol). Pure ethanol has a density of 0.789 g/mL. (water is the solvent and ethanol is the solute) What is the freezing point of the solution? (Kf (water)  1.86C/m)
1
Expert's answer
2020-02-27T09:04:55-0500

1. kH=c/p

kH=13.7/1000 x 1.5 = 6.326e-4

2. ∆Т= К x 1000 x m1/ m2 x M

∆T= 1.86 x 10.7 x 1000 / 125 x 1.44 = 110